Operations Report – April 27th

Crew 265 Operations Report Sol 3 27-04-2022


Name of person filing report: Dave Laude/Sergii Iakymov

Non-nominal systems: Furnace, Suit #8. Percy’s left rear tail light is non-functional, but the brake light is working.

Notes on non-nominal systems: Furnace is functional and we remain cautious over limiting to early morning only. Infrequently used now. Suit #8 has blown a fuse during our mission while on the table.

Spirit rover used: Yes

Hours: 174.3
Beginning charge: 100
Ending charge: 100!
Currently charging: Yes

Opportunity rover used: No
Beginning charge:
Ending charge:
Currently charging:

Curiosity rover used: No
Beginning charge:
Ending charge:
Currently charging:

Perseverance rover used: Yes
Hours: 236.5
Beginning charge: 100
Ending charge: 89
Currently charging: Yes

General notes and comments:

Summary of Hab operations:

Water (static tank): Not full. ~410 gallons.
Water (loft tank): 30 gallons
Water Meter: 158117.6 units

Static to Loft Pump used – Yes
Static tank pipe heater (on or off): Off
Static tank heater (On or off): Off
Toilet tank emptied: Yes

Summary of internet: Nominal

Summary of suits and radios: Batteries charging.
Suite #8 has been repaired (blown fuse has been replaced; loose wires soldered and isolated with electrical tape ). Additionally, we discovered a loose charging port in the same suite. After some adjustments and repairs the suit is functional again. The receiving connectors on the suit were slightly expanded open by repetitive insertions and removals of the charger plug making the charging dependent upon adjusting the plug to just the correct position for proper contact. Now just plug in and charge without fussing with it, for now. There are two other one-piece suits that require attention that we plan to get at soon.

Summary of GreenHab operations: NA

WATER USE: gallons

Heater: On or Off

Supplemental light: On or Off


Summary of ScienceDome operations: NA

Dual split: Heat or AC, On or Off

Summary of RAM operations: Used soldering iron, solder and tape to repair suit #8.

Summary of any observatory issues: NA

Summary of health and safety issues: Per HSO report one person is prohibited from EVA due to high blood pressure and another is being monitored for bp and heart rate a few times a day.

Questions, concerns and requests to Mission Support: (note trash pickup)

Getting hot in hab upper deck. What is the status of the swamp cooler water supply and any idea how much water it could consume?

Copyright © The Mars Society. All rights reserved. | Main Site